OpenStudy (anonymous):

Which equation is a quadratic equation?   3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3)

OpenStudy (anonymous):

Is the question which one is NOT a quadratic equatoin ? only the 1st one isn't. The 1st one is a linear equation.

OpenStudy (anonymous):

It's which one is a quadratic equation... How do you figure out if it is one?

OpenStudy (anonymous):

A quadratic equation will be equivalent to this: y=ax^2+bx+c The highest power of x will be 2 ( x^2 )

OpenStudy (anonymous):

Oh, okay... Thanks! :) I'll try and figure this one out then... :P

OpenStudy (anonymous):

wait the last one isnt quadratic

OpenStudy (anonymous):

oh, ok. :)

OpenStudy (anonymous):

Is this it ? y − 3 = (x^2 − 3)(8x + 3) If so: y = 8x^3 + 3x^2 - 24x -9 +3 y = 8x^3 + 3x^2 - 24x -6 The highest power of x above is 3 , so it's not a quadratic equation

OpenStudy (anonymous):

are you supposed to pick only one correct answer , or can you pick more than one answer ?

OpenStudy (anonymous):

Only one. :)

OpenStudy (anonymous):

make sure all the options in the question are written out correctly then. can you click the "Draw" tool below, and write the options there ?